OpenStudy (anonymous):

Kathy bicycles 4 km/hr faster than Carlos. In the same time it takes Carlos to bicycle 27 km, Kathy can bicycle 39 km. How fast does each bicyclist travel?

OpenStudy (anonymous):

use \[t=\frac{d}{r}\] and since the time is the same put \[\frac{27}{r}=\frac{39}{r+4}\] \[39r=27(r+4)\] \[39r=27r+156\] \[12r=156\] \[r=13\]

OpenStudy (anonymous):

carlos at 13, kathy at 17

OpenStudy (anonymous):

great thanks